Please Help..
Thanks...
Thanks...
You do not have the required permissions to view the files attached to this post.
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
ksh- I think that's because there are 4 different ways that that the "winning" scenario that Jim explained can occur (4 envelopes)ksh wrote:Hi jim,
The last part is not clear. why have u multiplied with 4. You have already taken into account 4 envelops and putting letters in 3 of them wrongly !
onesome..I have used a generic formula!! If u are unaware of it, better not follow that approach.onesome wrote:Hi getneonow
Im sorry but i could not understand your solution completely.
To be precise this part:
:
Now we have to see that the remaining 3 letters are put in the wrong envelopes which is given by (3!) [1/2! - 1/3! ]
--------------------
This is how i am trying to solve:
Letter 1 => Can be any one of the four letters , i.e. 4 ways to pick a letter - and 1 correct way to place that one into the right envelope.
Therefore, for the first letter => 4*1 ways to place in correct envelope
For the remaining 3 letter:
3 ways to pick aletter => 2 ways to place in incorrect envelope ==> 3*2
2 ways to pick a letter => 1 ways to place in incorrect envelope ==> 2*1
1 ways to pick a letter => 2 ways to place in incorrect envelope ==> 1*1
Total number of ways for (1 correct & 3 incorrect) = (4*1) * (3*2) * (2*1) * (1*1) = 48 ways
Total ways to place 4 letters in 4 envelopes => (4*4) * (3*3) * (2*2) * (1*1) = 576 ways
Probability (1 correct & 3 incorrect) = 48/576 = 1/12
----------------------------------------------------------------
I would appreciate if someone can explain how is this wrong.
And how to reach the correct solution.
Thanking in advance.
Note that you have already marked the letters here and specifically asking the first letter A to go to correct envelope.durgesh79 wrote:Could you please explain the approach if i make following change in the question.
The letters are marked A, B, C and D, Find out the probability the A goes in the right enevelope.
Bingo!!durgesh79 wrote:Thanks Neo,
Makes perfect sense.
Now the probability that only one of A, B, C or D goes into the right envelope will be 1/12+1/12+1/12+1/12 = 1/3